OpenStudy (chycora):

7. (3 pt) Solve. 4x – 6 = −3x + 8 x =

OpenStudy (chycora):

is this 15

OpenStudy (chycora):

@iGreen and @TreyJones31658

OpenStudy (r_josh10):

Let's solve your equation step-by-step. 4x−6=−3x+8 Step 1: Add 3x to both sides. 4x−6+3x=−3x+8+3x 7x−6=8 Step 2: Add 6 to both sides. 7x−6+6=8+6 7x=14 Step 3: Divide both sides by 7. 7x 7 = 14 7 x=2 Answer: x=2

OpenStudy (chycora):

k thanx